Circular No. 6048

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN SCORPIUS J. F. Schachter, Smithsonian Astrophysical Observatory, reports: “We

note that virtually the entire quoted BATSE error circle (IAUC 6046) lies in the Scorpius OB1 association (Perry et al. 1991, A.Ap. Suppl. 90, 195). A known Einstein Slew Survey x-ray source not associated with any known O or B star, 1ES 1649–403 (Elvis et al. 1992, Ap.J. Suppl. 80, 257), lies just at the northwest corner of Scorpius OB1 and is only 0^{o}.47 from the quoted BATSE position. The epoch for our detection was that of the Einstein IPC (1978–1981). The position of the 1ES source is α = 16^{h}49^{m}07^{s}, δ = −40^{o}20′.4 (equinox 1950.0; 90-percent-confidence error radius of 1′.2). Its unabsorbed flux in the band 0.2–4 keV, assuming a 5-keV bremsstrahlung spectrum and column density of ∼ 10^{22} cm^{−2}, is (1.80 ± 0.44) × 10^{−11} erg cm^{−2} s^{−1}. Assuming a distance of 2 kpc, the estimated quiescent luminosity is ∼ 10^{34} erg s^{−1}. Archival ROSAT PSPC data (during 1991 Feb.– Mar.) show a strong source near the position of 1ES 1649–403 (although the image’s proximity to an opaque support structure makes a precise determination of the offset difficult). The two evident ROSAT sources in the BATSE error circle are associated with early-type stars. We suggest that 1ES 1649–403 could be the BATSE source seen in quiescence. Imaging observations at Cerro Tololo are planned in the next few days.”

SUPERNOVA 1994W IN NGC 4041 C. M. Gaskell, Department of Physics and Astronomy, University of

Nebraska, writes: “The descriptions of the spectrum of SN 1994W (IAUC 6046) closely match the appearance of SN 1984E in spectra taken by me at maximum light. The narrow hydrogen lines in SN 1984E disappeared within a few weeks, presumably when the blast wave overtook the circumstellar shell (Henry and Branch 1987, PASP 99, 112; Gaskell and Keel 1988, in Supernova 1987A in the Large Magellanic Cloud, Cambridge, p. 13). Close monitoring of SN 1994W would be worthwhile to see if its narrow lines suffer a similar fate.”

Further visual magnitude estimates (cf. IAUC 6044): Aug. 3.90 UT, 13.6 (G. Cortini and M. Villi, Mt. Colombo, Italy); 7.23, 13.5 (C. E. Spratt, Victoria, BC, Canada).

1994 August 8 Daniel W. E. Green